OpenStudy (anonymous):

In a school fair there is a tombala stool There are 100 tickets numbered 1 to 100. The tickets with numbers ending in 5 or 0 win a prize. Lee picks two tickets at random without replacement. Find the probability that both of Lee's tickets win a prize show working out please

OpenStudy (anonymous):

\[(20÷100)×(19÷99)=19÷495=0.0383838\]

OpenStudy (anonymous):

5 and 0 ending numbers will get a token,, 100/5=20 >>20/100 >> 1/5 then, 100/10=10 >> 10/100 >> 1/10 the probability of winning is 1/5 x 1/10 = 1/50 Find the probability that both of Lee's tickets win a prize.. (given 2 tickets) 1/50 x 2 = 1/25; i hope i did it. I'm not that good in probability!!

OpenStudy (anonymous):

(20÷100)×(19÷99)=19÷495=0.0383838 Because when Lee is drawing first ticket, there are 20 winning number in 100 cards. After that when Lee is drawing second ticket, there are 19 winning number in 99 cards. We are assuming that Lee drawn a winning number at the first time otherwise we don't need to continua solving. So, answer is ; 0.0383838 (We can round it to the "%04".)

OpenStudy (anonymous):

hear2x cerkinner. mine was only in a fraction form.